A 150 mL solution of alcohol in water is 75% alcohol. How many mL of a 60% solution should be added to make the solution 72% alcohol?
Answer by TimothyLamb(4379)

150(0.75) + v(0.60) = (150 + v)(0.72)
112.5 + v(0.60) = 108 + v(0.72)
v(0.12) = 4.5
v = 4.5/(0.12)
v = 37.5
---
answer:
add 37.5 mL of 60% solution
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php